# Non-simple abelian varieties in a family and countable intersections on moduli space

### v1

@@ 1. Introduction
let $\mathcal{A}_{\_}{g}$ be the moduli space of principally polarized abelian varieties of dimension $g$. Let $K$ be a number field and Let $X\in\mathcal{A}_{\_}g/K$ be a rational curve parameterizing a one-dimensional family of abelian varieties. This is equivalent to saying $X\simeq\mathbb{P}^{1}$. Let $A_{\_}{x}$ denote the fiber of this family over $x\in X$ and let $A_{\_}\eta$ be the generic fiber. Assume that the geometric monodromy of $A_{\_}\eta$ modulo $\ell$ is the full symplectic group $\operatorname{Sp}_{\_}2g(\mathbb{F}_{\_}\ell)$ for almost all primes $\ell$. In particular, this implies that $A_{\_}\eta$ is geometrically simple. We are interested in understanding how this property extends to other fibers. To be precise, we ask the following question:
Question 1. What can we say about the number of $x\in X(\mathbb{Q})$ of bounded height such that $A_{\_}{x}$ is geometrically non-simple?
@@
One may ask whether $S(B)$ is finite. If so, can we obtain an upper bound in terms of $B$? The finiteness of $S(B)$ was proven in [EEHK09, Thm. A] under certain assumptions ensuring large monodromy for the family. Specifically, these assumptions concern a particular type of family of abelian varieties, namely the family
@@
However, the constant $C$ depends on the largest prime dividing the discriminant of $f$. In this article, we aim to generalize these results to arbitrary rational one-dimensional families of abelian varieties over a number field, without knowing the explicit defining equation of the family. Notably, compared to [EEHK09, Theorem B], the logarithmic terms in our theorem do not depend on $g$ or $K$. To state our theorem, we need the following definitions.
@@
Let $n$ be the transcendence degree of $\mathcal{A}_{g}$ over $K$ and let $N=n+1$. Define $\iota:X\to\mathbb{P}^{N}$ as the restriction of the Hodge embedding of $\mathcal{A}_{g}$ to $X$ (see also Section 2.1 for details).
Therefore, specifying $iota$ is the same as specifying the family $X$.
In the definition of $S(B)$, $H(t)$ is the height of $t$ as an element of $K$. Let $P_{\_}t=\iota(t).$ We now introduce the following definition.
@@
Let $K$ be a number field of degree $d_{\_}K$ over $\mathbb{Q}$. Let $X\in\mathcal{A}_{g}/K$ be a rational curve parameterizing a one-dimensional family of abelian varieties. Let $A_{\_}\eta$ be the generic fiber of this family, and assume that the geometric monodromy of $A_{\_}\eta$ modulo $\ell$ is the full symplectic group $\operatorname{Sp}_{\_}{2g}(\mathbb{F}_{\_}\ell)$ for almost all primes $\ell$. Let $\iota$ be the embedding defined in 1.2 and let $d$ denote the degree of $X$ with respect to $\iota$. Then there exists an absolute constant $C_{\_}{K,g}$, depending on $K$ and $g$, such that
@@
If we change the definition of the height from $H(P_{\_}t)$ to $H(t)$ and assume that $H(t)\leq B$, then there exists a constant $C_{\_}{\iota}$ depends on $d_{\_}K$, $g$ and the family $X$, and an absolute constant $\kappa$ such that
@@ 2.4. Parabolic subgroups of the symplectic groups
Let $G=\operatorname{Sp}(\psi)$ be the symplectic group for a symplectic space $(V,\psi)$ of dimension $2n$, where $n$ is a positive integer and $k$ is a field. We call an increasing flag $F={F^{i}}_{\_}i$ of nonzero proper subspaces of $V$ isotropic if each $F^{i}$ is isotropic with respect to the symplectic form $\psi$. Each parabolic $k$-subgroup of $G$ is the stabilizer of some isotropic flag. In particular, a maximal parabolic subgroup of $G$ is exactly the stabilizer of a minimal non-empty isotropic flag; i.e., the $G$-stabilizer of a non-zero isotropic subspace $F^{1}\subset V$. It is a standard fact that there are $n$ such conjugacy classes, corresponding to a choice of vertex to remove from the Dynkin diagram and to the dimension of $F^{1}$, which lies between $1$ and $n$.
@@ 3.1. The congruent covers
We study level structures on $\mathcal{A}_{\_}g$, the moduli space of principally polarized abelian varieties, formed by certain congruence subgroups $\Gamma$ of $\Gamma(1):=\mathrm{Sp}_{\_}{2g}(\mathbb{Z})$. We are particularly interested in the standard level subgroups $\Gamma(\ell)$, which encode principally polarized abelian varieties $A$ equipped with a symplectic basis of $A[\ell]$. This level structure is denoted by $\mathcal{A}{g,\Gamma(\ell)}$. The cover $\mathcal{A}_{\_}{g,\Gamma(\ell)}$ carries an action of $\mathrm{GSp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z}).$
Let $H$ be a subgroup of $\mathrm{GSp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z})$ that contains a maximal parabolic subgroup of $\mathrm{Sp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z}).$ There are several ways to represent the congruence quotient $\mathcal{A}_{\_}{g,\Gamma(\ell)}/H$: either via the cover $q:\mathcal{A}{g,\Gamma(\ell)}/H\to\mathcal{A}_{\_}g$, or via the modular correspondence in $\mathcal{A}_{\_}g\times\mathcal{A}g$, where a point on $\mathcal{A}{g,\Gamma(\ell)}/H$ can be identified with a pair $(A,F^{1})$ with
@@
Our goal is to bound the number of points with bounded height of $X(K)$, such that the corresponding moduli points are non-simple. Let $\mathcal{A}_{\_}x$ be the fiber over $x\in X(K)$. If $\mathcal{A}_{\_}x$ is non-simple, then the image of the Galois representation $\operatorname{Gal}_{\_}K\to\operatorname{GSp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z})$ is conjugate to some block matrices and therefore its intersection with $\operatorname{Sp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z})$ is contained in some maximal parabolic subgroups $H$. Let $\tilde{X}_{\_}H\subset\mathcal{A}_{\_}{g,\Gamma(\ell)}/H$ be the lift of $X$ and $\pi:\tilde{X}_{\_}H\to X$ be the restriction of $q$ to $\tilde{X}_{\_}H$. Our task now is to bound the number of points in $X(K)$ that lie in $\pi(\tilde{X}_{\_}H(K))$. The modular interpretation enables us to construct suitable projective embeddings to control the change of heights when lifting the base curve $X$. We will discuss these details in Section 4. We conclude this subsection with a summary in the following corollary:
@@
Every $k$-rational point $t\in S(B)$ lifts to some congruence cover $\mathcal{A}_{\_}{g,\Gamma(\ell)}/H$ for some prime integer $\ell$ and maximal parabolic subgroup $H\leq\operatorname{GSp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z})$.
@@ 3.2. The lift is integral
Let $U\subset X$ be the smooth locus of the family, and $\eta$ denote the generic point. Let $\ell$ be a prime integer. We say a family $\mathcal{A}$ of abelian varieties over $U$ has big monodromy modulo $\ell$ if the image of the geometric monodromy of $A_{\_}\eta$
@@
For any $g\geq 1$, there exists a constant _$\ell_{\_}1(g)\geq 1$ such that if _$\ell\geq\ell_{\_}1(g)$ and $A$ is an abelian variety of dimension $g$ over a field $k$ with big monodromy modulo $\ell$, then $\operatorname{End}(A)=\mathbb{Z}$ and, in particular, $A$ is geometrically simple ([EEHK09, Prop. 4]). However, the converse is not always true. For example, let $\mathcal{A}\to C$ be a family of abelian surfaces with real multiplication by $\mathbb{Q}(\sqrt{d})$. The monodromy group of $A_{\_}\eta$ modulo $\ell$ is either $\operatorname{GL}_{\_}2(\mathbb{F}_{\_}\ell)\times\operatorname{GL}_{\_}2(\mathbb{F}_{\_}\ell)$ or $\operatorname{GL}(\mathbb{F}_{\_}\ell^{2})$, yet $A_{\_}\eta$ can still be geometrically simple. To prove the integrality, we assume that the family $\mathcal{A}$ has big monodromy.
@@
The big monodromy condition is known for a large class of abelian varieties defined over finitely generated fields of arbitrary characteristic. The most prominent result of this type is the classical theorem of Serre (cf. [Ser84, 2.2.7], [Ser85, Lemme 1]) for the number field case. Its generalization to finitely generated fields of characteristic zero is well-known: If $A$ is an abelian variety over a finitely generated field $k$ of characteristic zero, with $\operatorname{End}_{\_}{\bar{k}}(A)=\mathbb{Z}$ and $\operatorname{dim}(A)=2,6$ or odd, then $A/k$ has big monodromy.
@@
In combination with the previous proposition we obtain the following corollary
@@ 4. Bounding change of heights
In this section, we provide an upper bound on the height of rational points in $\tilde{X}_{\_}{H}(K)$ lying over a point $P_{\_}t\in X(K)$, in terms of the height of $t$ and the level $\ell$. The main result of this section is Proposition 4.1.
Let $\iota_{\_}H^{\prime}$ denote the restriction of the product of Hodge embeddings to $\tilde{X}_{\_}{H}$ and let $\iota_{\_}H$ be the composition of $\iota_{\_}H^{\prime}$ with the Segre embedding. Corollary 3.1 asserts that, for a suitable choice of $\ell\in\mathbb{Z}$, there exist congruence covers of $X$ corresponding to maximal parabolic subgroups, such that a rational point $t\in X(K)$ with height at most $B$ lifts to a rational point on the cover. These covers can be arranged into the following commutative diagram:
@@
The constant $C_{\_}\iota$ is computable if the coefficient of $j_{\_}1,\dots,j_{\_}{n+1}$ are known.

### v2

# Non-simple abelian varieties in a family: arithmetic approaches

@@ 1. Introduction
let $\mathcal{A}_{\_}{g}$ be the moduli space of principally polarized abelian varieties of dimension $g$. Let $K$ be a number field and let $X\subset\mathcal{A}_{\_}g$ be a rational curve over $K$ parameterizing a one-dimensional family of abelian varieties. This is equivalent to saying $X$ is isomorphic to _$\mathbb{P}^{1}_{\_}{\bar{K}}$. For a $K$-point $x\in X$, let $A_{\_}{x}$ denote the specialization of this family over $x$. Let $A_{\_}\eta$ be the generic fiber of this family. Assume that the geometric monodromy of $A_{\_}\eta$ modulo $\ell$ is the full symplectic group $\operatorname{Sp}_{\_}{2g}(\mathbb{F}_{\_}\ell)$ for almost all primes $\ell$. In particular, this implies that $A_{\_}\eta$ is geometrically simple. We are interested in understanding how the property of being geometrically simple extends to other fibers. To be precise, we ask the following question:
Question 1. What can we say about the number of $x\in X(K)$ of bounded height such that $A_{\_}{x}$ is geometrically non-simple?
@@
One may ask whether $S(B)$ is finite. If it is finite, can we obtain an upper bound in terms of $B$? The finiteness of $S(B)$ was proven in [EEHK09, Thm. A] under certain assumptions that ensure large monodromy for the family. Specifically, these assumptions concern a particular type of family of abelian varieties, namely the family
@@
However, there is an error in their proof, which causes the constant $C$ depends on the largest prime dividing the discriminant of $f$. In this article, we aim to generalize these results to arbitrary rational one-dimensional families of abelian varieties over a number field, without knowing the explicit defining equation of the family. Notably, compared to [EEHK09, Theorem B], the logarithmic terms in our theorem do not depend on $g$ or $K$. To state our theorem, we need the following definitions.
@@
Let $n$ be the dimension of $\mathcal{A}_{g}$ over $K$ and let $N=n+1$. Define $\iota:X\to\mathbb{P}^{N}$ as the restriction of the Hodge embedding of $\mathcal{A}_{g}$ to $X$ (see also Section 2.1 for details).
We see that specifying $\iota$ is the same as specifying the family $X$.
In the definition of $S(B)$, $H(t)$ is defined as the height of $t\in X(K)$ as an element of $K$. Let $P_{\_}t=\iota(t)$. Alternatively, we can view $t$ as an element of $\mathbb{P}^{N}$ and define its height as the multiplicative height of $P_{\_}t$. Using this perspective, we introduce the notion of $S_{\_}X(B)$ as follows:
@@
Let $K$ be a number field of degree $d_{\_}K$ over $\mathbb{Q}$. Let $X\subset\mathcal{A}_{g}$ be a rational curve over $K$ parameterizing a one-dimensional family of abelian varieties. Let $A_{\_}\eta$ be the generic fiber of this family, and assume that the geometric monodromy of $A_{\_}\eta$ modulo $\ell$ is the full symplectic group $\operatorname{Sp}_{\_}{2g}(\mathbb{F}_{\_}\ell)$ for almost all primes $\ell$. Let $\iota$ be the embedding defined in 1.2 and let $d$ denote the degree of $X$ with respect to $\iota$. Then there exists an absolute constant $C_{\_}{K,g}$, depending on $K$ and $g$, such that
@@
If we change the height of $t$ from $H(P_{\_}t)$ to $H(t)$ and assume that $H(t)\leq B$, then there exists a constant $C_{\_}{\iota}$ depends on $d_{\_}K$, $g$ and the family $X$, and an absolute constant $\kappa$ such that
@@ 2.4. Parabolic subgroups of the symplectic groups
Let $G=\operatorname{Sp}(\psi)$ be the symplectic group for a symplectic space $(V,\psi)$ of dimension $2n$, where $n$ is a positive integer and $k$ is a field. We call an increasing flag $F=\{F^{i}\}_{\_}i$ of nonzero proper subspaces of $V$ isotropic if each $F^{i}$ is isotropic with respect to the symplectic form $\psi$. Each parabolic $k$-subgroup of $G$ is the stabilizer of some isotropic flag. In particular, a maximal parabolic subgroup of $G$ is exactly the stabilizer of a minimal non-empty isotropic flag; i.e., the $G$-stabilizer of a non-zero isotropic subspace $F^{1}\subset V$. It is a standard fact that there are $n$ such conjugacy classes, corresponding to a choice of vertex to remove from the Dynkin diagram and to the dimension of $F^{1}$, which lies between $1$ and $n$.
@@ 3.1. The congruent covers
We study level structures on $\mathcal{A}_{\_}g$, the moduli space of principally polarized abelian varieties, formed by certain congruence subgroups $\Gamma$ of $\Gamma(1):=\operatorname{Sp}_{\_}{2g}(\mathbb{Z})$. We are particularly interested in the standard level subgroups $\Gamma(\ell)$, which encode principally polarized abelian varieties $A$ equipped with a symplectic basis of $A[\ell]$. This level structure is denoted by $\mathcal{A}_{\_}{g,\Gamma(\ell)}$. The cover $\mathcal{A}_{\_}{g,\Gamma(\ell)}$ carries an action of $\operatorname{GSp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z}).$
Let $H$ be a subgroup of $\operatorname{GSp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z})$ that contains a maximal parabolic subgroup of $\operatorname{Sp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z}).$ There are several ways to represent the congruence quotient $\mathcal{A}_{\_}{g,\Gamma(\ell)}/H$: either via the cover $q:\mathcal{A}_{\_}{g,\Gamma(\ell)}/H\to\mathcal{A}_{\_}g$, or via the modular correspondence in $\mathcal{A}_{\_}g\times\mathcal{A}_{\_}{g}$, where a point on $\mathcal{A}_{\_}{g,\Gamma(\ell)}/H$ can be identified with a pair $(A,F^{1})$ with
@@
Our goal is to bound the number of points with bounded height on $X(K)$, such that the corresponding moduli points are non-simple. Let $\mathcal{A}_{\_}x$ be the fiber over $x\in X(K)$. If $\mathcal{A}_{\_}x$ is non-simple, then the image of the Galois representation $\operatorname{Gal}_{\_}{\overline{K}/K}\to\operatorname{GSp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z})$ is conjugate to some block matrices and therefore its intersection with $\operatorname{Sp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z})$ is contained in some maximal parabolic subgroups $H$. Let $\tilde{X}_{\_}H\subset\mathcal{A}_{\_}{g,\Gamma(\ell)}/H$ be the lift of $X$ and $\pi:\tilde{X}_{\_}H\to X$ be the restriction of $q$ to $\tilde{X}_{\_}H$. Our task now is to bound the number of points in $X(K)$ that lie in $\pi(\tilde{X}_{\_}H(K))$. Modular interpretation enables us to construct suitable projective embeddings to control height change when lifting the base curve $X$. We will discuss these details in Section 4. We conclude this subsection with a summary in the following corollary:
@@
Every $k$-rational point $t\in S(B)$ lifts to a congruence cover $\mathcal{A}_{\_}{g,\Gamma(\ell)}/H$ for some prime integer $\ell$ and a maximal parabolic subgroup $H\leq\operatorname{GSp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z})$.
@@ 3.2. The lift is integral
Let $U\subset X$ be the smooth locus of the family and $\eta$ denote the generic point. Let $\ell$ be a prime integer. We say a family $\mathcal{A}$ of abelian varieties over $U$ has big monodromy modulo $\ell$ if the image of the geometric monodromy of $A_{\_}\eta$
@@
For any $g\geq 1$, there exists a constant _$\ell_{\_}1(g)\geq 1$ such that if _$\ell\geq\ell_{\_}1(g)$ and $A$ is an abelian variety of dimension $g$ over a field $k$ with big monodromy modulo $\ell$, then $\operatorname{End}(A)=\mathbb{Z}$ and, in particular, $A$ is geometrically simple ([EEHK09, Prop. 4]). However, the converse is not always true. For example, let $\mathcal{A}\to C$ be a family of abelian surfaces with real multiplication by $\mathbb{Q}(\sqrt{d})$. The monodromy group of $A_{\_}\eta$ modulo $\ell$ is either $\operatorname{GL}_{\_}2(\mathbb{F}_{\_}\ell)\times\operatorname{GL}_{\_}2(\mathbb{F}_{\_}\ell)$ or $\operatorname{GL}_{\_}{2}(\mathbb{F}_{\_}\ell^{2})$, yet $A_{\_}\eta$ can still be geometrically simple. To prove the integrality, we assume that the family $\mathcal{A}$ has big monodromy.
@@
The big monodromy condition is known for a large class of abelian varieties defined over finitely generated fields of arbitrary characteristic. The most prominent result of this type is the classical theorem of Serre (cf. [Ser84, 2.2.7], [Ser85, Lemme 1]) for the number field case. Its generalization to finitely generated fields of characteristic zero is well-known: If $A$ is an abelian variety over a finitely generated field $k$ of characteristic zero, with $\operatorname{End}_{\_}{\overline{k}}(A)=\mathbb{Z}$ and $\operatorname{dim}(A)=2,6$ or odd, then $A/k$ has big monodromy.
@@
In combination with the previous proposition, we obtain the following corollary:
@@ 4. Bounding change of heights
In this section, we prove an upper bound on the height of rational points in $\tilde{X}_{\_}{H}(K)$ lying over a point $P_{\_}t\in X(K)$, in terms of the height of $t$ and the level $\ell$. The main result of this section is Proposition 4.1.
Let $\iota_{\_}H^{\prime}$ denote the restriction of the product of the Hodge embeddings to $\tilde{X}_{\_}{H}$ and let $\iota_{\_}H$ be the composition of $\iota_{\_}H^{\prime}$ with the Segre embedding. Corollary 3.1 says that, for a suitable choice of $\ell\in\mathbb{Z}$, there exist congruence covers of $X$ corresponding to maximal parabolic subgroups such that a rational point $t\in X(K)$ with height at most $B$ elevates to a rational point on the cover. These covers can be arranged into the following commutative diagram:
@@
The constant $C_{\_}\iota$ is computable if the coefficients of $j_{\_}1,\dots,j_{\_}{n+1}$ are known.